# Explaining Explainability: Understanding Concept Activation Vectors

### v1

@@ 6.1.1 Theory
$f({\bm{a}}_{l_{1}}+{\bm{v}}_{c,l_{1}})$ $=f({\bm{a}}_{l_{1}})+{\bm{v}}_{c,l_{2}}.$
Assume we have found ${\bm{v}}_{c,l_{1}}$ and would like to find a ${\bm{v}}_{c,l_{2}}$ that satisfies Eq. 11. If $f$ conserves vector addition, such as for a linear layer, then it holds that:
$f({\bm{a}}_{l_{1}})+f({\bm{v}}_{c,l_{1}})$ $=f({\bm{a}}_{l_{1}})+{\bm{v}}_{c,l_{2}}$ (12)
${\bm{v}}_{c,l_{2}}$ $=f({\bm{v}}_{c,l_{1}}).$
Hence, it is possible to have consistent vectors across layers if $f$ conserves vector addition and ${\bm{v}}_{c,l_{2}}=f({\bm{v}}_{c,l_{1}})$. Alternatively, if $f$ does not conserve vector addition, we cannot simplify Eq. 11 and for a given ${\bm{v}}_{c,l_{1}}$:
${\bm{v}}_{c,l_{2}}=f({\bm{a}}_{l_{1}}+{\bm{v}}_{c,l_{1}})-f({\bm{a}}_{l_{1}}).$ (13)
If ${\bm{v}}_{c,l_{2}}$ depends on ${\bm{a}}_{l_{1}}$, there does not exist a ${\bm{v}}_{c,l_{2}}$ such that Eq. 11 is true for all ${\bm{a}}_{l_{1}}$. In other words, there does not exist a vector in layer $l_{2}$ which has the same effect on the activations as the vector in layer $l_{1}$ for all inputs to the model.
ReLU and sigmoid functions are common NN activations. Using Eq. 13, for any input $i$, if
$f=\text{ReLU}:\quad a_{l_{1},i}+v_{l_{1},i}>0,\ a_{l_{1},i}\leq 0,\text{ or }a_{l_{1},i}+v_{l_{1},i}\leq 0,\ a_{l_{1},i}>0$ (14)
$f=\text{sigmoid}:\quad v_{l_{1},i}\neq 0$ (15)
then there does not exist a consistent ${\bm{v}}_{c,l_{2}}$, i.e. it is impossible to have consistent vectors under these conditions. The proofs for Eqs. 14 and 15 are available in Appendix 10. Next, we demonstrate that we cannot find consistent vectors in practice.
6.1.2 Experiments
Our goal is to investigate the question are the concept vectors found using TCAV consistent? We measure the consistency of two perturbations using the consistency error:
@@ 6.1.2 Experiments
Optimised CAV
(lower bound): TCAV may not find a ${\bm{v}}_{c,l_{2}}$ that has a consistency error of 00 with ${\bm{v}}_{c,l_{1}}$. Therefore, we use gradient decent on ${\bm{v}}_{c,l_{2}}$ to minimise the consistency error, which acts as a lower bound.
Projected CAV
: the error between $f({\bm{v}}_{c,l_{1}})$ and ${\bm{v}}_{c,l_{2}}$, which measures how consistent the vectors are when projected into the next layer. If $f(\cdot)$ conserves vector addition, the projected CAVs would have 00 error.
Random
(upper bound): We include two benchmarks. Random CAVs found using probe datasets containing random images, and a Random Direction vector: ${\bm{v}}_{c,l_{2}}\sim\text{Uniform}(-1,1)$. If the consistency error is similar to random, it suggests that the CAVs between layers are as similar to each other as random directions.
Fig. 3 shows the $\epsilon_{consistency}$ for different ${\bm{v}}_{c,l_{2}}$ across different training runs (see Appendix 13 for details). The concept CAV obtains a nonzero consistency error, suggesting that CAVs across different layers are not consistent. When we compare it with the benchmarks, we find:
  • The consistency error for the optimised CAVs is lower, implying that the standard approach to find CAVs does not find optimally consistent CAVs. However, the nonzero error for optimised CAVs suggests it is not possible to find consistent vectors in these layers.
  • The projected CAVs have a nonzero error, indicating that vector addition is not preserved.
@@
Figure 3: Empirical evidence for inconsistent CAVs. The consistency error for different ${\bm{v}}_{c,l_{2}}$ for striped in the penultimate convolutional layer of a ResNet-50 trained on ImageNet. The optimised CAV acts as lower bound, whereas the random CAV and Direction act as baselines that provide an inutitive upper bounds. Concept CAV: striped CAVs, trained as normal. Projected CAV: striped CAVs from layer $l_{1}$ projected into layer $l_{2}$, $f({\bm{v}}_{c,l_{1}})$.
The inability to find consistent concept vectors across layers suggests that the directions encoded by CAVs in different layers are not equivalent; instead we speculate that they represent different components of the same concept. This aligns with the intuition that model representations are more complex later in the NN [30, 32, 5], therefore it is unlikely that the same aspects of a concept are represented in different layers (discussed further in Appendix 13.3). Consequentially, TCAV scores across layers can vary as they perform different tests – they measure the class sensitivity to a different version of the concept.
Fig. 5(c) shows that concept vectors found in different layers of a model can give contradictory TCAV scores (further examples available in Appendix 13.4). In the Elements dataset, shape concepts are encoded in each layer as the test accuracy for each layer is above $93\%$. Therefore, we expect to be able to use TCAV on each of these layers. However, the TCAV scores for cross in the Elements dataset contradict each other across ‘layers.3’ and ‘layers.4’, suggesting a positive and negative influence, respectively. This contradiction makes it difficult to draw a conclusion about the model’s class sensitivity to cross.
@@ 6.2 Entanglement
Next, we investigate the effect of entangled concept vectors on TCAV score. We analyse the TCAV scores for the ‘striped triangles’ class in $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$. The class label depends solely on the presence stripes and triangle. Therefore, we expect all other concepts to obtain low TCAV scores (indicating negative sensitivity), as their presence makes the class less likely, or insignificant TCAV scores, if the concept is uninformative.222assuming that the model uses each concept correctly
@@ 6.3 Spatial Dependence
Finally, we investigate NH3: are CAVs spatially dependent? We reshape the CAVs back into the original shape of the activations, and compute the channel-wise norm as follows:
@@
When a spatially independent probe dataset is used to create CAVs, as in the top row of Fig. 6, the spatial norms are uniform, suggesting the CAVs are not spatially dependent333the individual CAVs may still be spatially dependent, but this cancels out across training runs. See Appendix 15.3 for details.. However, when the probe dataset exhibits spatial dependence, so do the resulting CAVs. The regions of near-zero norm indicate that the corresponding spatial regions of the gradients do not contribute to the directional derivative and, consequently, to the TCAV score.
Next, we investigate the question does the model have a different conceptual sensitivity depending on the concepts location in the input image? As CAVs operate in the activation space of a specific layer, we can show that model is not translation invariant if:
@@
    That each depth-wise slice of the activations, of shape $(1,1,D)$, affects the logit output differently.
@@
To address (2), we compute the TCAV scores for different sets of spatially dependent CAVs to determine if the sensitivity of the model changes depending on the concepts location. To investigate this, we created spatially dependent classes in the Elements dataset, where the class depends on what concepts are present and on where they are in the image, such as ‘striped triangles on the left’. We use spatially dependent CAVs to show that a model is not translation invariant with respect to striped or triangle in Fig. 5(b). Here, we discuss the results for the class of ‘striped triangles on the left’. The TCAV scores for striped, triangle, striped left and triangle left are high, indicating a positive influence of these concepts on the class. However, the striped right and triangle right TCAV scores often do not differ significantly from the null scores, providing no evidence to suggest the model is sensitive to these concepts. The difference between the left and right biased TCAV scores indicates that the model is not translation invariant with respect to these concepts as the model’s sensitivity depends on where the concept is present in the image input space. Overall, this suggests that we can use CAVs to detect model translation invariance. See Appendix 15.6 for examples on ImageNet.
@@ 7 Practitioner Recommendations
In Section 5, we provided example papers which use CAVs and may be influenced by the above properties. As a more concrete example, we examine the use-case of Yan et al. [42] which uses CAVs in the context of skin cancer diagnosis. Some of the concepts have expected spatial dependencies, for example, dark borders and dark corners. Spatial norms could be used to confirm these spatial dependencies exist. Equally, for concepts such as the presence of a ruler, the spatial norms could confirm the CAVs have no overall spatial dependence. There are multiple concepts which have opposed meanings, for example regular streaks and irregular streaks. The cosine similarities between CAVs could confirm that these concepts are negatively correlated or at least less similar to each other than to other concepts. We provide a more detailed analysis of this use-case in Appendix 17.

### v2

@@ 6.1 Consistent CAVs
$f({\bm{a}}_{l_{1}}+{\bm{v}}_{c,l_{1}})$ $=f({\bm{a}}_{l_{1}})+{\bm{v}}_{c,l_{2}}.$ (12)
Assuming $f$ is continuous and differentiable, in Appendix 9 we prove the result that Eq. 12 can hold if and only if $f$ is equal to
$f({\bm{a}}_{l_{1}})=g({\bm{a}}_{l_{1}})+M{\bm{a}}_{l_{1}}+{\bm{b}}.$ (13)
Where $g(\cdot)$ is a periodic function with period ${\bm{v}}_{c,l_{1}}$ and $M\in\mathbb{R}^{m_{l_{2}}\times m_{l_{1}}}$ is a non-zero linear term with constant ${\bm{b}}\in\mathbb{R}^{m_{l_{2}}}$. More intuitively, we can obtain layer consistent vectors ${\bm{v}}_{c,l_{1}}$ and ${\bm{v}}_{c,l_{2}}$ if and only if $f$ is composed of a periodic function with period ${\bm{v}}_{c,l_{1}}$ and a non-zero linear term $M$. In principal, $f$ could approximate a function of this form since neural networks are universal approximators (Sonoda & Murata, 2017). However, it seems reasonably unlikely that even if the model was modeling a periodic function it would have a period of exactly the same direction as a CAV. Throughout the rest of this section we provide empirical evidence that, in practice, layer consistent CAVs are not found.
Experiments
Our goal is to investigate the question are the concept vectors found using TCAV consistent across layers? We measure the consistency of two perturbations using the consistency error:
@@
Optimised CAV (lower bound): TCAV may not find a ${\bm{v}}_{c,l_{2}}$ that has a consistency error of 00 with ${\bm{v}}_{c,l_{1}}$. Therefore, we use gradient decent on ${\bm{v}}_{c,l_{2}}$ to minimise the consistency error, which acts as a lower bound.
Projected CAV: the error between $f({\bm{v}}_{c,l_{1}})$ and ${\bm{v}}_{c,l_{2}}$, which measures how consistent the vectors are when projected into the next layer. If $f(\cdot)$ conserves vector addition, the projected CAVs would have 00 error.
Random (upper bound): We include two benchmarks. Random CAVs found using probe datasets containing random images, and a Random Direction vector: ${\bm{v}}_{c,l_{2}}\sim\text{Uniform}(-1,1)$. If the consistency error is similar to random, it suggests that the CAVs between layers are as similar to each other as random directions.
Figure 3 shows the $\epsilon_{consistency}$ for different ${\bm{v}}_{c,l_{2}}$ across different training runs (see Appendix 12 for details). The concept CAV obtains a nonzero consistency error, suggesting that CAVs across different layers are not consistent. When we compare it with the benchmarks, we find:
  • The consistency error for the optimised CAVs is lower, implying that the standard approach to find CAVs does not find optimally layer consistent CAVs. However, the nonzero error for optimised CAVs suggests it is not possible to find consistent vectors across these layers.
  • As expected, the projected CAVs have a nonzero error, indicating that vector addition is not preserved.
@@
Figure 3: Empirical evidence for inconsistent CAVs across layers. The consistency error for different ${\bm{v}}_{c,l_{2}}$ for striped in the penultimate convolutional layer of a ResNet-50 trained on ImageNet. The optimised CAV acts as lower bound, whereas the random CAV and Direction act as baselines that provide an intuitive upper bounds. Concept CAV: striped CAVs, trained as normal. Projected CAV: striped CAVs from layer $l_{1}$ projected into layer $l_{2}$, $f({\bm{v}}_{c,l_{1}})$.
The inability to find consistent concept vectors across layers suggests that the directions encoded by CAVs in different layers are not equivalent; instead we speculate that they represent different components of the same concept. This result in unsurprising given that previous works have demonstrated that model representations are more complex later in the NN (Mordvintsev et al., 2015; Olah et al., 2017; Bau et al., 2017), therefore it is unlikely that the same aspects of a concept are represented in different layers (discussed further in Appendix 12.3). Consequentially, TCAV scores across layers can vary as they perform different tests – they measure the class sensitivity to a different version of the concept.
Figure 5(c) shows that concept vectors found in different layers of a model can give contradictory TCAV scores (further examples available in Appendix 12.4). In the Elements dataset, shape concepts are encoded in each layer as the test accuracy for each layer is above $93\%$. Therefore, we expect to be able to use TCAV on each of these layers. However, the TCAV scores for cross in the Elements dataset contradict each other across ‘layers.3’ and ‘layers.4’, suggesting a positive and negative influence, respectively. This contradiction makes it difficult to draw a conclusion about the model’s class sensitivity to cross.
@@
In order to determine how often TCAV scores are inconsistent across layers, we introduce a new metric, the TCAV layer consistency score. It is a measure of how well the significant TCAV scores for some concept, $c$, and class, $k$, agree with each other across different layers $l_{i},i\in 1\dots L$. It is based on how many layers are on the same side of the null TCAV score. The null TCAV score is the mean TCAV score for a set of random CAVs (see Appendix 10.1 for more detail). Mathematically the metric, $S_{\text{consistent}}$, is defined as
$S_{\text{consistent}}=|2((\frac{1}{L}\sum_{i}\operatorname{TCAV}_{c,k,l_{i}}>\operatorname{TCAV}_{r,k,l_{i}})-0.5)|$ (15)
where $r$ indicates the TCAV score is calculated using random CAVs. It is difficult to make confident statements about the sensitivity of the model when the layer consistency score is close to 00 as it indicates the TCAV scores from different layers tend to disagree on the direction of the sensitivity.
We obtain a mean TCAV consistency score of $0.841$ across all concepts/classes/layers for Elements and a mean TCAV consistency score of $0.868$ for a selection of $6$ classes and $16$ concepts for a ResNet-50 trained on ImageNet (See Appendix 12.5 for further details). The scores indicate that although complete disagreement across layers is not common ($5.5\%$ and $2.1\%$ of scores were equal to 00 for Elements and ImageNet, respectively), it can occur, and a not insubstantial number of concepts/classes have a reasonable amount of disagreement across layers ($23\%$ and $14\%$ of scores were $\leq 0.5$ for Elements and ImageNet, respectively).
@@ 6.2 Entanglement
Next, we investigate the effect of entangled concept vectors on TCAV scores. We analyse the TCAV scores for the ‘striped triangles’ class in $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$. The class label depends solely on the presence stripes and triangle. Therefore, we expect all other concepts to obtain low TCAV scores (indicating negative sensitivity), as their presence makes the class less likely, or insignificant TCAV scores, if the concept is uninformative.333assuming that the model uses each concept correctly
@@ 6.3 Spatial Dependence
Finally, we investigate NH3: are CAVs spatially dependent? In the case of a convolutional based neural network, where the activations are of shape $H\times W\times D$, we can reshape the CAVs back into the original shape of the activations, and compute the channel-wise norm as follows:
@@
When a spatially independent probe dataset is used to create CAVs, as in the top row of fig. 6, the spatial norms are uniform, suggesting the CAVs are not spatially dependent444the individual CAVs may still be spatially dependent, but this cancels out across training runs. See Appendix 14.3 for details.. However, when the probe dataset exhibits spatial dependence, so do the resulting CAVs. The regions of near-zero norm indicate that the corresponding spatial regions of the gradients do not contribute to the directional derivative and, consequently, to the TCAV score.
Next, we investigate the question does the model have a different conceptual sensitivity depending on the concept’s location in the input image? As CAVs operate in the activation space of a specific layer, we can show that a model is not translation invariant if:
@@
    Each depth-wise slice of the activations, of shape $(1,1,D)$, affects the logit output differently.
@@
To address (2), we compute the TCAV scores for different sets of spatially dependent CAVs to determine if the sensitivity of the model changes depending on the concepts location. To investigate this, we created spatially dependent classes in the Elements dataset, where the class depends on what concepts are present and on where they are in the image, such as ‘striped triangles on the left’. We use spatially dependent CAVs to show that a model is not translation invariant with respect to striped or triangle in fig. 5(b). Here, we discuss the results for the class of ‘striped triangles on the left’. The TCAV scores for striped, triangle, striped left and triangle left are high, indicating a positive influence of these concepts on the class. However, the striped right and triangle right TCAV scores often do not differ significantly from the null scores, providing no evidence to suggest the model is sensitive to these concepts. The difference between the left and right biased TCAV scores indicates that the model is not translation invariant with respect to these concepts as the model’s sensitivity depends on where the concept is present in the image input space. Overall, this suggests that we can use CAVs to detect model translation invariance. See Appendix 14.6 for examples on ImageNet and in Appendix 14.7, even though we cannot use spatial norms to visualise it, some preliminary evidence that spatially dependent CAVs can exist for transformer-based architectures.
@@ 7 Practitioner Recommendations
In § 5, we provided example papers do not consider these properties but may be influenced by them. To provide a concrete example, we examine the use-case of Yan et al. (2023) which uses CAVs in the context of skin cancer diagnosis. Below we demonstrate how our recommendations could have been used and how the analysis impacts the conclusions drawn.
Consistency
Entanglement
Spatial Dependence
7.1 Experiment Setup
To further this example, we run illustrative experiments on a similar dataset to Yan et al. (2023). The dataset used in Yan et al. (2023) is not publicly available.
Model
We finetune a ResNet50 (He et al., 2016) pretrained on ImageNet (Deng et al., 2009) on the ISIC $2019$ dataset (Tschandl et al., 2018; Codella et al., 2017; Combalia et al., 2019) for the binary classification of melanoma. We use a binary cross entropy loss and the Adam optimiser (Kingma & Ba, 2015), training until convergence of validation loss to achieve an area under the receiver operating characteristic curve (AUC) of $0.91$ on the validation split.
CAVs
For the CAVs, as in Yan et al. (2023), we use the derm7pt dataset (Kawahara et al., 2019). There are $12$ clinical concepts which have been expertly labelled within the dataset. We hand labelled three additional concepts, which were used in Yan et al. (2023), of dark corners, dark borders and ruler which are possible confounders for the model. We defined dark corners as any image with a circular aperture which left the corners of the image black, dark borders as any image containing rectangles of blacked out areas and ruler as any image containing a ruler. For each of the medical concepts, there are three labels: typical/regular, atypical/irregular, and absent. When training CAVs for these concepts, we either used random images or images with the label of ‘absent’ as the negative probe dataset. We used the ‘absent’ labels because using random images as the negative set gave low-quality CAVs, with accuracies of $50$-$65$% (see fig. 7). It is unclear from Yan et al. (2023) what they used for negative sets. For training the CAVs we used $70$ images per concept and used $30$ different random seeds for the random negative probe set to get $30$ CAVs per concept. Yan et al. (2023) do not use TCAV, so this setup is different from the original paper.
7.2 Results
As shown in fig. 7, initial results with random CAVs gave poor performance for the medical concepts, so we used the ‘absent’ category for each class as the negative set. This gave better performances with accuracy between $60$-$75\%$ for the medical concepts but this is still far lower than for the confounders at $80$-$90\%$. We hypothesise that this is due to the simplicity of the confounding concepts. This is supported by the accuracy reported by Yan et al. (2023), where they also obtained a lower CAV accuracy for the medical concepts. The accuracy for the confounding concepts of dark borders and dark corners drops in later layers. This is likely due to the spatial nature of the concepts – an idea further supported by fig. 10 where the CAVs in later layers have reduced spatial distinction.
Below, we examine how the three CAV properties analysed in this paper affect the TCAV scores and the conclusions you can draw from them.
Figure 7: Mean CAV test accuracies for the melanoma use-case. Top: Medical concepts where random images (right) or images where the concept is labelled as absent (left) are used in the negative probe dataset. Bottom: Potential confounders where CAVs were trained with (right) and without (left) a flip augmentation.
Consistency
The TCAV scores for many of the medical concepts (fig. 8(a)) are consistent across layers, irrespective of if random images or ‘absent’ images were used for the negative probe set. The consistent scores provide more confidence in using them to explain the model’s behaviour as repeated significance tests are performed indicating the model has the same sensitivity to the concept. In terms of understanding the model, the scores provide some evidence that it operates similar to human experts, as the TCAV scores for the atypical/irregular medical concepts are high for the malignant class, as expected, and the confounding concepts are often not significant (prior to using a flip augmentation - see the spatial dependence section below for discussion), providing little evidence that the model is sensitive to the confounders.
(a) TCAV scores are not qualitatively different for CAVs of differing accuracy. TCAV scores for medical concepts where random images (left) or images where the concept is labelled as absent (right) are used in the negative probe dataset.
(b) CAVs trained with an augmentation were significant more often. TCAV scores for potential confounders where CAVs were trained with and without a flip augmentation.
Figure 8: TCAV scores for the melanoma use-case. The standard deviation is black or red for significant and insignificant results, respectively. The null for each layer is shown as a horizontal black line.
Entanglement
Interestingly, for the CAVs with ‘absent’ labelled images in their negative probe dataset (the right of fig. 9), the CAVs with opposed meanings often appear to be the most similar to each other. For example, regular vascular structures has a negative or zero similarity with all concepts except irregular vascular structures (with a similarity of $0.11$). We hypothesise that this is because the two concepts share the same negative set. This hypothesis is supported by Ramaswamy et al. (2022a) where, while they did not discuss changing solely the negative set, they did show that CAVs are sensitive to the choice of probe dataset. In addition, if compared to the similarities between CAVs trained using random images as the negative set (the left of fig. 9), we see that this pattern disappears. The higher similarity between concepts which have opposite meanings suggests that the CAVs do not represent the concepts they are labelled for. Therefore, in this case, we do not believe the TCAV scores can be interpreted as we do not have confidence the CAVs represent their desired concept. This example highlights a more general problem that the negative probe set can have a substantial affect on the resulting CAV, even though it is the positive probe set that is designed to represent the concept.
Figure 9: Cosine similarity matrix for CAVs of different concepts from derm7pt when random images (left) or images where the concept is labelled as absent (right) are used in the negative probe dataset.
Figure 10: Mean CAV spatial norms for a selection of CAVs from the melanoma use-case.
Spatial Dependence
The spatial norms in fig. 10 show a clear spatial dependence in the center for each of the medical concepts across all layers. This aligns with expectations, as the dataset requires the skin lesion to be centered in the image and each concept is related to the appearance of the lesion. The dark corners and dark borders concepts, however, show deviations to this pattern in the earlier layers, with dark corners having high spatial norms in the corners and dark borders high spatial norms in the top. This is desirable, as these confounding concepts depend on features at the edge of the image, away from the lesion. Along with the accuracies in fig. 7, this suggests that the CAVs in earlier layers better represent these spatially dependent concepts. For dark borders, however, we hypothesise that the high spatial norms in just a single direction suggest that the CAVs are not a good representation of dark borders in locations other than the top. Therefore, we retrained CAVs for each of the confounding concepts but with an augmentation applied to the probe dataset to randomly flip the images in the horizontal and/or vertical direction. This removes any bias that there may be in the probe dataset for the concept to be in one particular direction. Figure 10 shows that the CAVs trained for dark borders with an augmentation (dark borders flip) had only minor improvements for their spatial norms, with some layers being slightly less uni-directional. Figure 7, however, shows that the accuracy for each of the CAVs improved. This suggests that there was an issue with bias in the probe dataset and by using a flip augmentation we create CAVs which better represent the concepts, but, surprisingly, the activations of the model encode the dark borders concept in a spatially unsymmetrical manner.
The TCAV scores for CAVs trained with the augmentation (fig. 8(b)) obtain significance in more layers but the results are still fairly inconclusive. For example, the TCAV scores for ruler in layer3.4 and layer 4.2 are above/below the null, respectively. This means that layer3.4 suggests a positive influence, whereas layer4.2 suggests a negative influence. With less than a $1\%$ difference in accuracy between the CAVs of the two layers it is not clear which layer we might trust more and so no conclusive statements can be said about the influence of ruler on malignancy predictions.
7.3 Summary
The poor accuracy of medical concepts when using random images in the negative probe set required the use of the ‘absent’ category instead. However, the similarity matrices in fig. 9 suggest that the CAVs trained using the ‘absent’ category do not represent the desired concepts. Therefore, we do not think either set of medical CAVs can produce meaningful TCAV scores.
For the confounders, however, the CAVs had high accuracy and the spatial norms indicated spatial dependence in concepts we expect spatial dependence, providing evidence that these CAVs are suitable to use. However, the spatial dependence seemed directional, and so an augmentation was added to flip the probe images horizontally/vertically. Although this did not substantially change the spatial norms, it improved CAV accuracy and increased the number of layers for which we had significant TCAV scores. From these scores, it appears the model is sensitive to dark corners although the evidence is weak with large standard deviations in TCAV score and for ruler we found inconclusive results with inconsistent scores across layers.
This use-case demonstrates the importance of analysing consistency, entanglement and spatial dependence of CAVs, alongside more typical evaluations such as CAV accuracy and statistical significance, in order to understand CAV-based explanations and the conclusions you can draw from them. Our experiments provide an example for practitioners to follow in their own experiments with CAVs.